HIGH-RESOLUTION MICROWAVE AND INFRARED MOLECULAR-BEAM STUDIES OF THE CONFORMERS OF 1,1,2,2-TETRAFLUOROETHANE

S.C. Stone, L.A. Philips, Department of Chemistry, Cornell University, Ithaca, NY 14853; \underline{G.T. Fraser}, F.J. Lovas, L.-H. Xu, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899; S.W. Sharpe, Molecular Sciences Research Center, Pacific Northwest Laboratory, Richland, WA 99352.

High-resolution microwave and infrared molecular-beam spectra have been measured for 1,1,2,2-tetrafluorethane (HFC134). For the higher-energy, polar, C_{1}-symmetry, gauche conformer, microwave spectra have been recorded for the normal and mono-^{13}C isotopomers and analyzed to determine a C-C bond length of 1.512(4) Å and a CCF angle of 109.7(3)^{∘}. A tunable microwave-sideband CO_{2} laser and electric-resonance optothermal spectrometer have been used to measure the infrared spectrum of the ν_{6}, C-C stretch of the gauche conformer near 906 cm^{−1}. Microwave-infrared double resonance and precise ground- state combination-differences provided by the microwave measurements guide the assignment of the spectrum. A b- and c-type spectrum is observed and fit to a Watson asymmetric-top Hamiltonian to within the experimental uncertainty of 0.3 MHz. The high quality of the fit and the similarity of the centrifugal distortion constants to the ground-state values indicate that the band is effectively unperturbed. A number of strong unassigned lines are present in the spectrum. These transitions do not display any microwave-infrared double resonance effect. The attribution of these transitions to the nonpolar anti conformer is ruled out since the transition intensities are sensitive to the field strength of the inhomogeneous electric field used to focus the molecules. Pulsed slit-jet diode-laser spectra have been recorded for the ν_{16}, anti conformer near 1127 cm^{−1}. An a- and c-type hybrid band is observed consistent with previous low-resolution assignments of this vibration to a B_{u} mode. A total of 522 non-blended transitions were assigned and fit to determine ground- and excited-state constants. The ground- state constants of A = 5134.952(65), B = 3148.277(27), and C = 2067.106(43) MHz are the first experimental determination of rotational constants for this conformer.